OpenStudy (anonymous):

Factor completely: 64a2 – 4b2

OpenStudy (anonymous):

Do you mean 64a² - 4b²? If yes, then: = 4(16a²-b²) = 4 (4a - b) (4a + b)

OpenStudy (anonymous):

yes i got it right thx

OpenStudy (anonymous):

(8a+2b)(8a-2b) or 2(4a+b)(4a-b)